Maths

 Here find the range and the coefficient of range for the following value 25,32,85,32,42,10,20,18,28
Answer #1

Range =H-L=85-10=75

Coefficient of range =H-L/H+L=85-10/85+10=75/95=.79

answered by: Vijayan